OpenStudy (anonymous):

What are the steps for solving a Maclaurin series for f(x): ln((1+x)/(1-x))

OpenStudy (anonymous):

i think i would start with \[\ln(1+x)-\ln(1-x)\]

OpenStudy (anonymous):

series for these are well known, although i guess you should derive them

OpenStudy (anonymous):

we could do them by hand, or use a trick either way works

OpenStudy (anonymous):

so then for example f'(x) = -2/(x^2-1), f"(x)=4x/(x^2-1)^2, etc, then don't you insert a value determined by x_o then simplify?

OpenStudy (anonymous):

whoa way way too much

OpenStudy (anonymous):

it will get very ugly very fast doing it that way break it apart first

OpenStudy (anonymous):

that is, start with \[\ln(1+x)-\ln(1-x)\] compute separately and subtract, it will be way simpler

OpenStudy (anonymous):

roger that, ultimately i'm trying to show that the original equation is equal to 2(x+(x^3/3)+(x^5/5)+(x^(2k-1)/(2k-1)+...), I probably should have clarified that in the beginning

OpenStudy (anonymous):

here is the very simplest way put \[f(x)=\ln(1+x)\]and then \[f'(x)=\frac{1}{1+x}\]

OpenStudy (anonymous):

\[\frac{1}{1+x}\] expands as a geometric series, namely \[1-x+x^2-x^3+x^4-...\] take the antiderivative term by term to get the expansion for \[\ln(1+x)\]

OpenStudy (anonymous):

alright that makes more sense

OpenStudy (anonymous):

you get \[\ln(1+x)=x-\frac{x^2}{2}+\frac{x^3}{3}-\frac{x^4}{4}+...\]

OpenStudy (anonymous):

this one is not even bad to do by hand \[f(x)=\ln(1+x)\] \[f'(x)=(1+x)^{-1}\] \[f''(x)=-(1+x)^{-2}\] \[f^{(3)}(x)=2(1+x)^{-3}\] and so on then replace \(x\) by 0 and divide by \(k!\) to the get expansion, but the first method is slicker

OpenStudy (anonymous):

repeat the process with \[\ln(1-x)\] take the derivative, recognize the geometric series, then integrate term by term it is quick

OpenStudy (anonymous):

ahhh alright, that makes a lot more sense

OpenStudy (anonymous):

have fun

OpenStudy (anonymous):

thanks a lot